OpenStudy (anonymous):

determine the zeros of the polynomial function : f(x) = x^4 - 13x^2 +36

OpenStudy (ksaimouli):

factor it set =0

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

x^4 - 13x^2 + 36 = 0

OpenStudy (anonymous):

then factor

OpenStudy (anonymous):

http://www.purplemath.com/modules/factquad.htm

OpenStudy (anonymous):

Set x^2 = t, and now you have a quadratic in t which you can solve by the factoring method. Once you have your t's, convert back to x.

OpenStudy (campbell_st):

ok just look at the problem this way \[(x^2)^2 - 13(x^2) + 36 = 0\] think of the x^2 as a single value then factorise \[(x^2 - 4)(x^2 - 9) = 0\] now solve for x

OpenStudy (anonymous):

(x-3)(X+3)(x-2)(x+2) ?

OpenStudy (campbell_st):

thats correct so you will have 4 solutions

OpenStudy (anonymous):

cant you just write it like (x^2-4)(x^2-9)? < in factored form

OpenStudy (anonymous):

It is not completely factored that way. If you look at each factor it still has an x^2 in it so you have to go further for the 4 solutions.

OpenStudy (anonymous):

okay makes sense, thanks!

OpenStudy (campbell_st):

no as you need to find the zeros, this can only be done by solving, and this polynomial can be solved by factoring